Find a c-inverse of the matrix
A D
2
4
1 0
2
1
Expert's answer
2020-04-09T15:10:04-0400

"AD=\begin{bmatrix} 
2 & 4 \\
10 & 2 \\
\end{bmatrix}"


inverse is given by,

"AD^{-1}=\frac{1}{4-40}\begin{bmatrix} 
2 & -4 \\
-10 & 2 \\
\end{bmatrix}"


"AD^{-1}=\begin{bmatrix} 
-0.0556 & 0.1111 \\
0.2778 & -0.0556 \\
\end{bmatrix}"